Solve using square roots (x+3)2 = 36


1
Expert's answer
2021-02-24T14:09:32-0500

"(x+3)^2=36"

"|x+3|=6"

"x+3=6" or "x+3=-6"

"x=3" "x=-9"

Answer: x=3 and x=-9